# Direct Detection of Dark Photon Dark Matter with the James Webb Space Telescope

### v2

@@ Acknowledgement
JL would like to thank Le Hoang Nguyen for the helpful discussions. The work of HA is supported in part by the National Key R&\&&D Program of China under Grant No. 2021YFC2203100 and 2017YFA0402204, and the NSFC under Grant No. 12475107. The work of SG is supported by NSFC under Grant No. 12247147, the International Postdoctoral Exchange Fellowship Program, and the Boya Postdoctoral Fellowship of Peking University. The work of JL is supported by NSFC under Grant No. 12075005, 12235001.
@@
The next-to-leading order term is suppressed by $e^{-2z^{2}}$, and can be neglected. In addition, note that that
@@
×$\times e^{-\frac{1}{8}k_{0}^{2}|\bm{r^{\prime}}-\bm{r^{\prime\prime}}|^{2}}\mathrm{Re}\left(\frac{e^{im_{A^{\prime}}(|\bm{r}-\bm{r^{\prime}}|-|\bm{r}-\bm{r^{\prime\prime}}|)}}{|\bm{r}-\bm{r^{\prime}}|^{3}|\bm{r}-\bm{r^{\prime\prime}}|^{2}}\right)$ (A.5)
@@
in the large $\alpha$ limit. In this analysis, we assume that $f(x)$ is second order differentiable, $g(x)$ is continuous. We also assume that $g(x)$ is either compactly supported or exhibits exponential decay and that $f^{\prime}(x)=0$ has only a discrete set of solutions. Here, we will focus on proving the case where $g(x)$ is compactly supported, noting that a similar approach can be applied when $g(x)$ has exponential decay.
@@
where $\delta$ is a positive real number that satisfies $\delta<\min\{\Delta,\alpha^{-1/2+\epsilon}\}$, with $\epsilon$ being a real number in the range $80<\epsilon<1/$ 8. Due to the continuity of $f^{\prime}(x)$, it follows that $f(x)$ is monotonic between any two adjacent points in $\Sigma$. Consequently, we can divide the interval $I$ into a finite set of closed intervals, each of which is either an element of $A$ or an interval on which $f(x)$ is strictly monotonic.
@@
where $c^{\prime}=f(c)$, $d^{\prime}=f(d)$, and $\tilde{g}(y)=g(f^{-1}(y))(f^{-1})^{\prime}(y)$. It can be shown that this expression goes to zero as $\alpha\rightarrow\infty$. By dividing the interval $[c^{\prime},d^{\prime}]$ into a set of intervals with length $2\pi/\alpha$, integrating on each small interval contributes a result of order $\mathcal{O}(\alpha^{-3/2})$. Summing over all intervals gives an additional factor of $\alpha$, resulting in the overall order $\mathcal{O}(\alpha^{-1/2})$. Consequently, (B.5) vanishes in the large $\alpha$ limit.
Next consider the integral on elements of $A$:
@@
It is straight forward to generalize to higher dimensional case. We can use the same method to prove that
@@
To evaluate the multidimensional Gaussian integral, we diagonalize the Hessian matrix. This transforms the integral into the product of $n$ one dimensional Gaussian integrals, ultimately leading to equation (2.6).
@@ Appendix C Ray Transfer Matrix Analysis
In systems satisfying paraxial condition, we can utilize ”ray transfer matrix analysis” to simplify the calculations. A beam of light can be characterized by two parameters: the angle(counterclockwise) between the light and the optical axis, and the vertical distance(upward) between the light and the optical axis. These two parameters can be organized into a column vector
@@
Here we assume that the light is travelling from left to right.
@@
[…]
@@ Appendix D Data Analysis Method
Here, we consider the parameter $a$’s as nuisance parameters. $S$ represents the DPDM-induced signal, and we assume its location to be in bin $i_{0}$. It’s worth noting that the frequency dispersion of DPDM is $\mathcal{B_{\rm DPDM}}\sim 0.15{\leavevmode\nobreak\ \rm kHz}\times(m_{A^{\prime}}/\mu{\rm eV})$. This is much smaller than the instrumental spectral resolution which ranges from 10 GHz to 40 THz, depending on different observation modes [56], so the DPDM-induced signal can be safely confined within a single frequency bin.

### v3

@@ Acknowledgement
JL would like to thank Le Hoang Nguyen for the helpful discussions. The work of HA is supported in part by the National Key R&\&D Program of China under Grant No. 2021YFC2203100 and 2017YFA0402204, and the NSFC under Grant No. 12475107. The work of SG is supported by NSFC under Grant No. 12247147, the International Postdoctoral Exchange Fellowship Program, and the Boya Postdoctoral Fellowship of Peking University. The work of JL is supported by NSFC under Grant No. 12075005, 12235001.
@@
The next-to-leading order term is suppressed by $e^{-2z^{2}}$, and can be neglected. In addition, note that
@@
$\times e^{-\frac{1}{8}k_{0}^{2}|\bm{r^{\prime}}-\bm{r^{\prime\prime}}|^{2}}\mathrm{Re}\left(\frac{e^{im_{A^{\prime}}(|\bm{r}-\bm{r^{\prime}}|-|\bm{r}-\bm{r^{\prime\prime}}|)}}{|\bm{r}-\bm{r^{\prime}}|^{3}|\bm{r}-\bm{r^{\prime\prime}}|^{2}}\right)$ (A.5)
@@
in the large $\alpha$ limit. In this analysis, we assume that $f(x)$ is second-order differentiable, $g(x)$ is continuous. We also assume that $g(x)$ is either compactly supported or exhibits exponential decay and that $f^{\prime}(x)=0$ has only a discrete set of solutions. Here, we will focus on proving the case where $g(x)$ is compactly supported, noting that a similar approach can be applied when $g(x)$ has exponential decay.
@@
where $\delta$ is a positive real number that satisfies $\delta<\min\{\Delta,\alpha^{-1/2+\epsilon}\}$, with $\epsilon$ being a real number in the range $0<\epsilon<1/8$. Due to the continuity of $f^{\prime}(x)$, it follows that $f(x)$ is monotonic between any two adjacent points in $\Sigma$. Consequently, we can divide the interval $I$ into a finite set of closed intervals, each of which is either an element of $A$ or an interval on which $f(x)$ is strictly monotonic.
@@
where $c^{\prime}=f(c)$, $d^{\prime}=f(d)$, and $\tilde{g}(y)=g(f^{-1}(y))(f^{-1})^{\prime}(y)$. It can be shown that this expression goes to zero as $\alpha\rightarrow\infty$. By dividing the interval $c^{\prime},d^{\prime}]$ into a set of intervals with length $2\pi/\alpha$, integrating on each small interval contributes a result of order $\mathcal{O}(\alpha^{-3/2})$. Summing over all intervals gives an additional factor of $\alpha$, resulting in the overall order $\mathcal{O}(\alpha^{-1/2})$. Consequently, (B.5) vanishes in the large $\alpha$ limit.
Next, consider the integral on elements of $A$:
@@
It is straightforward to generalize to higher-dimensional case. We can use the same method to prove that
@@
To evaluate the multidimensional Gaussian integral, we diagonalize the Hessian matrix. This transforms the integral into the product of $n$ one-dimensional Gaussian integrals, ultimately leading to equation (2.6).
@@ Appendix C Ray Transfer Matrix Analysis
In systems satisfying the paraxial condition, we can utilize ”ray transfer matrix analysis” to simplify the calculations. A beam of light can be characterized by two parameters: the angle(counterclockwise) between the light and the optical axis, and the vertical distance(upward) between the light and the optical axis. These two parameters can be organized into a column vector
@@
Here, we assume that the light is travelling from left to right.
@@
$\begin{bmatrix}1&-\frac{2}{\rho}\\ 0&1\end{bmatrix}\mbox{(concave)}\qquad\begin{bmatrix}1&\frac{2}{\rho}\\ 0&1\end{bmatrix}\mbox{(convex)}$ (C.3)
@@ Appendix D Data Analysis Method
Here, we consider the parameter $a$’s as nuisance parameters. $S$ represents the DPDM-induced signal, and we assume its location to be in bin $i_{0}$. It’s worth noting that the frequency dispersion of DPDM is $\mathcal{B_{\rm DPDM}}\sim 0.15{\penalty 10000\ \rm kHz}\times(m_{A^{\prime}}/\mu{\rm eV})$. This is much smaller than the instrumental spectral resolution which ranges from 10 GHz to 40 THz, depending on different observation modes [36], so the DPDM-induced signal can be safely confined within a single frequency bin.